### v1

@@ 1 Introduction
That being said, the field of causal discovery is predominantly method-driven rather than application-driven: the community produces new methods and algorithms at high speed but still relies on toy datasets and simple metrics for their evaluation (Gentzel et al., 2019), impeding its development and applicability to real-world problems. Recently, a plethora of surveys of causal discovery have covered existing causal discovery methods (Wang et al., 2024; Hasan et al., 2023; Zanga et al., 2022; Assaad et al., 2022; Vowels et al., 2022; Zhou and Chen, 2022; Nogueira et al., 2021; Guo et al., 2020; Glymour and Zhang, 2019; Malinsky and Danks, 2018; Singh et al., 2017), but none focused on the datasets and real-world applications to which these methods were applied. However, using good datasets and benchmarks is just as crucial as having good algorithms. For example, this has been pivotal in the recent deep learning boom with datasets such as ImageNet (Deng et al., 2009) and its associated challenge (Russakovsky et al., 2015). Beyond the choice of datasets, there is also a need for deeper consideration of the types of problems to which causal discovery can and should be applied. Over-reliance on simple settings makes the field disconnected from real-world challenges, and without practical applications, causal discovery risks becoming merely theoretical storytelling.
@@ 2 Background
Causal discovery. The task of recovering the graph $G$ from a dataset $\mathcal{D}$ (possibly containing interventional data) is called causal discovery. Constraint-based methods such as the PC algorithm (Spirtes et al., 2001) perform conditional independence tests to recover $G$, while score-based methods achieve this by finding the graph that maximizes a score, such as the Bayesian Information Criterion (BIC) (Glymour and Zhang, 2019). Some hybrid methods combine aspects of both approaches. Other methods make parametric assumptions on the functional form of the causal mechanisms or the variable distributions and orient edges based on detected asymmetries. For a more complete presentation, see Vowels et al. (2022). All of these methods are only guaranteed to recover the correct “ground-truth” graph in the infinite sample limit if the data-generating mechanism satisfies specific assumptions.
@@
Even when all these assumptions are satisfied, causal discovery is a hard task, both combinatorically and statistically. The space of DAGs scales super-exponentially with respect to the number of variables, and the assumptions above only guarantee correctness in the infinite sample limit, while in practice, one also has to deal with finite sample errors. Moreover, when the data is purely observational, without further assumptions one can at best identify an equivalence class of graphs, called the Markov Equivalence Class. Utilizing interventions represents the optimal strategy for overcoming obstacles in causal identifiability since it can greatly shrink the size of the equivalence class. If single-target interventions are performed on every node except one, the ground-truth graph is identifiable and, in general, fewer interventions are required (Eberhardt et al., 2012).
@@ 3.1 Synthetic Datasets
Synthetic datasets are usually generated by following these steps: first, the causal graph is sampled, then the causal mechanisms parameters, and finally, the data is sampled using ancestral sampling (i.e., by sampling the variables following their topological ordering). Common approaches include the Erdős-Rényi scheme, which uniformly samples a DAG, and scale-free networks (Barabási and Albert, 1999), which have been promoted as being more realistic (Barabási, 2009). The causal mechanisms parameters often follow a particular functional form assumed by the causal discovery method. For instance, one of the most common causal mechanisms is linear relations with Gaussian noise. Alternatives include nonlinear (Peters et al., 2014) and post-nonlinear additive noise models (Zhang and Hyvarinen, 2012) which are often used since they lead to identifiability results.
However, as highlighted in Reisach et al. (2021) and Reisach et al. (2024), the way these datasets are generated can sometimes be problematic as it introduces artifacts that some causal discovery methods may exploit. Namely, Reisach et al. (2021) showed that one can recover the causal ordering of some synthetic datasets simply by sorting the variables according to their variances. Even when these pitfalls are avoided by changing the data generating process (Andrews and Kummerfeld, 2024; Ormaniec et al., 2024), synthetic datasets are much simpler than their real-world counterpart and thus, the performance of proposed causal discovery methods is overestimated (Eigenmann et al., 2020). As we will elaborate in Section 5.4, real-world problems rarely conform to many of the assumptions built into synthetic datasets. We finish by noting that synthetic datasets can still be used for more realistic evaluation by benchmarking causal discovery methods on synthetic datasets where these stringent assumptions are violated, as in Montagna et al. (2024).
@@ 4 Systematic Review of the Causal Discovery Literature
Fig 2 shows the distribution of dataset types used in the selected studies. We observe that 20% of these only make use of purely synthetic datasets, while 64% rely on real-world datasets. Most real-world datasets are small, with 80% containing 20 or fewer variables. Fig 3 shows the field of provenance of pseudo-real and real-world datasets. Biology is by far the most prevalent field. This is partly explained by the ubiquitous use of the flow cytometry dataset, often simply named Sachs. It is the only real-world dataset considered for 35% of all the papers relying on real-world datasets (we redo our analysis excluding it in Appendix A.4). For the pseudo-real datasets, the most commonly used datasets come from the bnlearn repository (Scutari, 2009; Friedman et al., 1997).
These two widely used datasets have some notable limitations. For the Sachs datasets, the consensus network used is not fully consistent with the one given in Sachs et al. (2005), in particular, the cycles are often omitted; the ground truth provided is not definitive (see Ramsey and Andrews (2018); Mooij et al. (2020)), and varies between studies. To address these issues, we provide an updated ground truth model in Appendix B.3. Unfortunately, the existence of a ground truth network leads to an over-reliance on structural metrics (less than 5% of studies use interventional metrics) - even though the dataset includes several interventions. Also, some studies rely on a pseudo-real version of the dataset generated by fitting a model to a consensus network (without cycles) (Scutari, 2009). Finally, the dataset is often not discriminatory for causal discovery methods: the reported performance has peaked at an SHD of around 12. For the pseudo-real datasets from bnlearn, most studies rely only on structural metrics and we note that the way the datasets are generated, all common assumptions are respected.
Table 1 summarizes the types of metrics used to evaluate performance on simulated (synthetic and pseudo-real) versus real-world datasets. Structural metrics dominate for simulated datasets, where the ground-truth graph is known (100% of studies use them). In contrast, evaluations on real-world datasets depend more heavily on qualitative assessments (36% vs. 3%). Structural metrics are also widely used for real-world datasets (67%) which can be explained by the overreliance on real-world datasets that contain a ground-truth graph. Both observational and interventional metrics are rare across dataset types, used in fewer than 10% of studies. Overall, most studies rely solely on structural metrics: 86% for the simulated data and 54% for the real-world data. We also note that for many real-world datasets that do contain interventions, interventional metrics were not used 89% of the time.
@@ 5.1 Biology: Biomolecular Networks
As mentioned in Section 4, the Sachs dataset (Sachs et al., 2005) is by far one of the most commonly used real-world datasets in the causal discovery community. Sachs et al. (2005) pioneered the use of individual cells as the smallest observational unit of intact biological systems, catapulting the available dataset size from dozens or hundreds (of mice, patients, etc) to thousands of cells. This is a flow cytometry dataset that includes abundance measurements for 11 proteins and phospholipids over 7466 CD4+ T cells exposed to nine perturbation conditions. Causal discovery algorithms are often applied to a limited version of the dataset that includes only the 5846 measurements from these seven conditions, see e.g., Wang et al. (2017), Yang et al. (2018), and Squires et al. (2020). The popularity of this dataset is partially accounted for by the fact that, in contrast with the papers above, Sachs et al. (2005) introduced a consensus network based on existing biological literature (which we update in Fig. 7). In the absence of such a network, one often resorts to comparing against partial ground truth, e.g., as done by Frot et al. (2019), who compare against the reference database TRRUST (Han et al., 2015).
@@ 5.2 Neuroscience
Neuroscience is often concerned with understanding mechanisms, which ultimately is about causality (Ross and Bassett, 2024). It distinguishes the connectome, which describes the wires – the observable physical connections between neurons or brain areas – from the effectome, which describes the causal influences between brain regions (Pospisil et al., 2024). And when it comes to causality, there is a wide spectrum of approaches, including those that assume that correlation is causation and those that ask for perturbations (Siddiqi et al., 2022). There has been growing interest in how we can uncover genuine causal relationships from neuronal recordings, establishing causal inference as a central paradigm in neuroscience research. (Reid et al., 2019).
@@ 6 Conclusion
Our conclusion in recommending the use of empirical datasets echoes the one from Gentzel et al. (2019). To be clear, synthetic datasets are useful, but they should be complemented by more realistic evaluations on pseudo-real and real-world datasets. When interventional data are present, good quantitative evaluation on real-world datasets exists. However, in many fields besides biology, interventional data are hard to come by and thus pseudo-real datasets might be more adequate. They conserve most of the synthetic datasets’ advantages while being more realistic. However, the creation of pseudo-real datasets should always remain grounded by considering real-world datasets and the assumptions they violate. Through this review, which compiles an extensive list of both simulators and empirical datasets, we aim to motivate researchers to diversify their dataset usage, moving beyond the confines of synthetic data to embrace the complexity and richness of the real world in their causal discovery endeavors.
@@
PB acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and acknowledges Assya Trofimov for helpful discussions. JW received funding from the European Research Council (ERC) Starting Grant CausalEarth under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 948112). CS received funding from Valence Labs and acknowledges Jason Hartford for helpful discussions. AD acknowledges Sara Magliacane for helpful discussions. KS was funded in part by NIMH grant 1R44MH135465. DS acknowledges support from NSERC Discovery Grant RGPIN-2023-04869, and a Canada-CIFAR AI Chair.
@@ Appendix B Biological data
As aforementioned, mRNA readout of gene expression is highly informative of cell state. However, one has to keep in mind that the relation between genes is always mediated: genes themselves do not execute functions in the cells, but rather the proteins which are created based on the information encoded in the genes. Roughly speaking, mRNA (from Gene X) translates into protein (to Protein X). However, factors such as RNA stability and degradation strongly affect this relationship. Hence, the Gene Regulatory Networks (GRN) being modeled via transcriptomics are actually carried out not by the quantified gene mRNAs, but by latent variables: the gene-encoded proteins. One way to model biomolecular networks more directly is via proteomic datasets, especially ones in which the abundance of the activated proteins is quantified (Sheng Wu and Radvanyi, 2010; Sachs et al., 2005; PTM, 2022), though these tend to be more challenging experimentally.
@@
B.3 The Sachs dataset
We present in Fig. 7 an updated version of the ground-truth graph of the Sachs dataset. We note that compared to most “consensus” networks used in causal discovery, this graph contains cycles. The cycle most firmly established and most likely to be detected is the feedback loop from Erk to Raf, which has multiple routes and may be either negative or positive, depending on the intermediate steps. We also note that since biology is variable and most of the edges are indirect, some may not be detectable. As highlighted in the main text, while we provide a ground-truth graph, we recommend not just using structural metrics, but also interventional metrics. We hope that researchers will consistently use this version.
Figure 7: Causal network as presented in Sachs et al. (2005), with additional edges since established. Added edges are in purple; note that all of these may introduce feedback cycles into the regulatory network.
@@ C.1 Links to datasets
In the following tables, we provide links to the different datasets we discussed in the main text. In Table 4, 5, 6 and 7 we report the links for pseudo-real, biology, neuroscience and Earth science datasets. We also keep an updated version of these lists of datasets at our github repo. The lists contains mainly datasets that have been used in causal discovery studies. Several are particularly interesting since they contain interventional data and can violate some common assumptions. Still, we advise thoughtful use in line with our recommendations, rather than applying them blindly.
@@
(Cheng et al., 2023) CausalTime link to dataset
@@
Note that the datasets from Replogle et al. (2022) have been processed and curated by (Chevalley et al., 2022). Also, Lopez et al. (2022) offer code to process the data from Frangieh et al. (2021).
@@
DANDI Large collection of modern large neuroscience datasets, including optogenetics (DANDI) raw data All data
(Dorkenwald et al., 2022) Drosophila connectome (Flywire) Connectome Simulated Ca2+ activities
@@
(Teeters and Sommer, 2009) Spiking data from various sources (CRCNS) database -
@@
(Shah et al., 2018) rs-fMRI data from the medial temporal lobe raw dataset -
(Poldrack et al., 2015) Hippocampal rs-fMRI (MyConnectome project) raw dataset -
@@
(Van Essen et al., 2013) rs-fMRI (Human Connectome Project) raw dataset -
@@
(Wang et al., 2003) Task fMRI (star/plus experiment) - -

### v2

@@ 1 Introduction
That being said, the field of causal discovery is predominantly method-driven rather than application-driven (Rolnick et al., 2024): the community produces new methods and algorithms at high speed but still relies on toy datasets and simple metrics for their evaluation (Gentzel et al., 2019), impeding its development and applicability to real-world problems. Recently, a plethora of surveys of causal discovery have covered existing causal discovery methods (Wang et al., 2024; Hasan et al., 2023; Zanga et al., 2022; Assaad et al., 2022; Vowels et al., 2022; Zhou and Chen, 2022; Nogueira et al., 2021; Guo et al., 2020; Glymour and Zhang, 2019; Malinsky and Danks, 2018; Singh et al., 2017), but none focused on the datasets and real-world applications to which these methods were applied. However, using good datasets and benchmarks is just as crucial as having good algorithms. For example, this has been pivotal in the recent deep learning boom with datasets such as ImageNet (Deng et al., 2009) and its associated challenge (Russakovsky et al., 2015). Beyond the choice of datasets, there is also a need for deeper consideration of the types of problems to which causal discovery can and should be applied. Over-reliance on simple settings makes the field disconnected from real-world challenges, and without practical applications, causal discovery risks becoming merely theoretical storytelling.
@@ 2 Background
Causal discovery. The task of recovering the graph $G$ from a dataset $\mathcal{D}$ (possibly containing interventional data) is called causal discovery. Constraint-based methods such as the PC algorithm (Spirtes et al., 2001) perform conditional independence tests to recover $G$, while score-based methods achieve this by finding the graph that maximizes a score, such as the Bayesian Information Criterion (BIC) (Chickering, 2002). Some hybrid methods combine aspects of both approaches. Other methods make parametric assumptions on the functional form of the causal mechanisms or the variable distributions and orient edges based on detected asymmetries. For a more complete presentation, see for example Glymour and Zhang (2019). All of these methods are only guaranteed to recover the correct “ground-truth” graph in the infinite sample limit if the data-generating mechanism satisfies specific assumptions.
@@
Even when all these assumptions are satisfied, causal discovery is a hard task, both combinatorically and statistically. The space of DAGs scales super-exponentially with respect to the number of variables, and the assumptions above only guarantee correctness in the infinite sample limit, while in practice, one also has to deal with finite sample errors. Moreover, when the data is purely observational, without further assumptions one can at best identify an equivalence class of graphs, called the Markov Equivalence Class. Utilizing interventions represents the optimal strategy for overcoming obstacles in causal identifiability since it can greatly shrink the size of the equivalence class. If single-target interventions are performed on every node except one, the ground-truth graph is identifiable and, in general, fewer interventions are required (Eberhardt et al., 2005).
@@ 3.1 Synthetic Datasets
Synthetic datasets are usually generated by following these steps: first, the causal graph is sampled, then the causal mechanisms parameters, and finally, the data is sampled using ancestral sampling (i.e., by sampling the variables following their topological ordering). Common approaches to sample a DAG include the Erdős-Rényi scheme, which samples directed edges with a fixed probability, and scale-free networks (Barabási and Albert, 1999), which samples edges following a preferential attachment process (Barabási, 2009). The causal mechanisms parameters often follow a particular functional form assumed by the causal discovery method. For instance, one of the most common causal mechanisms is linear relations with Gaussian noise. Alternatives include nonlinear (Peters et al., 2014) and post-nonlinear additive noise models (Zhang and Hyvarinen, 2012) which are often used since they lead to identifiability results.
However, as highlighted in Reisach et al. (2021) and Reisach et al. (2024), the way these datasets are generated can sometimes be problematic as it introduces artifacts that some causal discovery methods may exploit. Namely, Reisach et al. (2021) showed that one can recover the causal ordering of some synthetic datasets simply by sorting the variables according to their variances. Even when these pitfalls are avoided by changing the data generating process (Andrews and Kummerfeld, 2024; Ormaniec et al., 2024), synthetic datasets are much simpler than their real-world counterpart and thus, the performance of proposed causal discovery methods is overestimated (Eigenmann et al., 2020). As we will elaborate in Section 5.4, real-world problems rarely conform to many of the assumptions built into synthetic datasets. We finish by noting that synthetic datasets can still be used for more realistic evaluation by benchmarking causal discovery methods in scenarios where these stringent assumptions are violated, as in Montagna et al. (2024).
@@ 4 Systematic Review of the Causal Discovery Literature
Fig 2 shows the distribution of dataset types used in the selected studies. We observe that 20% of these only make use of purely synthetic datasets, while 62% rely on real-world datasets. Most real-world datasets are small, with 80% containing 20 or fewer variables. Fig 3 shows the field of provenance of pseudo-real and real-world datasets. Biology is by far the most prevalent field. This is partly explained by the ubiquitous use of the flow cytometry dataset, often simply named Sachs (Sachs et al., 2005). It is the only real-world dataset considered for 35% of all the papers relying on real-world datasets (we redo our analysis excluding it in Appendix A.4). For the pseudo-real datasets, the most commonly used datasets come from the bnlearn repository (Scutari, 2009; Friedman et al., 1997).
These two widely used datasets have some notable limitations. For the Sachs datasets, the consensus network used is not fully consistent with the one given in Sachs et al. (2005), in particular, the cycles are often omitted; the ground truth provided is not definitive (see Ramsey and Andrews (2018); Mooij et al. (2020)), and varies between studies. Unfortunately, the existence of a ground truth network leads to an over-reliance on structural metrics (less than 5% of studies use interventional metrics) - even though the dataset includes several interventions. Also, some studies rely on a pseudo-real version of the dataset generated by fitting a model to a consensus network (without cycles) (Scutari, 2009). Finally, the dataset is often not discriminatory for causal discovery methods: the reported performance has peaked at an SHD of around 12. For the pseudo-real datasets from bnlearn, most studies rely only on structural metrics and we note that the way the datasets are generated, all common assumptions are respected.
Table 1 summarizes the types of metrics used to evaluate performance on simulated (synthetic and pseudo-real) versus real-world datasets. Structural metrics dominate for simulated datasets, where the ground-truth graph is known (100% of studies use them). In contrast, evaluations on real-world datasets depend more heavily on qualitative assessments (36% vs. 3%). Structural metrics are also widely used for real-world datasets (67%) which can be explained by the overreliance on real-world datasets that contain a ground-truth graph. Both observational and interventional metrics are rare across dataset types, used in fewer than 10% of studies. Overall, most studies rely solely on structural metrics: 86% for the simulated data and 54% for the real-world data. We also note that for real-world datasets that do contain interventions, interventional metrics were not used 89% of the time.
@@ 5.1 Biology: Biomolecular Networks
As mentioned in Section 4, the Sachs dataset (Sachs et al., 2005) is by far one of the most commonly used real-world datasets in the causal discovery community. Sachs et al. (2005) pioneered the use of individual cells as the smallest observational unit of intact biological systems, catapulting the available dataset size from dozens or hundreds (of mice, patients, etc) to thousands of cells. This is a flow cytometry dataset that includes abundance measurements for 11 proteins and phospholipids over 7466 CD4+ T cells exposed to nine perturbation conditions. Causal discovery algorithms are often applied to a limited version of the dataset that includes only the 5846 measurements from these seven conditions, see e.g., Wang et al. (2017), Yang et al. (2018), and Squires et al. (2020). The popularity of this dataset is partially accounted for by the fact that, in contrast with the papers above, Sachs et al. (2005) introduced a consensus network based on existing biological literature. In the absence of such a network, one often resorts to comparing against partial ground truth, e.g., as done by Frot et al. (2019), who compare against the reference database TRRUST (Han et al., 2015).
@@ 5.2 Neuroscience
Neuroscience is often concerned with understanding mechanisms, which ultimately is about causality (Ross and Bassett, 2024). It distinguishes the connectome, which describes the wires – the observable physical connections between neurons or brain areas – from the effectome, which describes the causal influences between brain regions (Pospisil et al., 2024). And when it comes to causality, there is a wide spectrum of approaches, including those that assume that correlation is causation and those that ask for perturbations (Siddiqi et al., 2022). There has been growing interest in how we can uncover genuine causal relationships from neuronal recordings, establishing causal inference as a central paradigm in neuroscience research (Reid et al., 2019).
@@ 6 Conclusion
Our conclusion in recommending the use of empirical datasets echoes the one from Gentzel et al. (2019). To be clear, synthetic datasets are useful, but they should be complemented by more realistic evaluations on pseudo-real and real-world datasets. When interventional data are present, good quantitative evaluation on real-world datasets exists. However, in many fields besides biology, interventional data are hard to come by and thus pseudo-real datasets might be more adequate. They conserve most of the synthetic datasets’ advantages while being more realistic. Still, the creation of pseudo-real datasets should always remain grounded by considering real-world datasets and the assumptions they violate. Through this review, which compiles an extensive list of both simulators and empirical datasets, we aim to motivate researchers to diversify their dataset usage, moving beyond the confines of synthetic data to embrace the complexity and richness of the real world in their causal discovery endeavors. We hope this effort will also foster the collaboration between causal discovery researchers and practitioners, leading to even more relevant real-world datasets and better integration of domain knowledge.
@@
PB acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and acknowledges Assya Trofimov for helpful discussions. JW received funding from the European Research Council (ERC) Starting Grant CausalEarth (Grant Agreement No. 948112) and from the German Federal Ministry of Education and Research (BMBF) as part of the project MAC-MERLin (Grant Agreement No. 01IW24007). CS received funding from Valence Labs and acknowledges Jason Hartford for helpful discussions. AD acknowledges Sara Magliacane for helpful discussions. KS was funded in part by NIMH grant 1R44MH135465. DS acknowledges support from NSERC Discovery Grant RGPIN-2023-04869, and a Canada-CIFAR AI Chair.
@@ Appendix B Biological data
As aforementioned, mRNA readout of gene expression is highly informative of cell state. However, one has to keep in mind that the relation between genes is always mediated: genes themselves do not execute functions in the cells, but rather the proteins which are created based on the information encoded in the genes. Roughly speaking, mRNA (from Gene X) translates into protein (to Protein X). However, factors such as RNA stability and degradation strongly affect this relationship. Hence, the Gene Regulatory Networks (GRN) being modeled via transcriptomics are actually carried out not by the quantified gene mRNAs, but by latent variables: the gene-encoded proteins. One way to model biomolecular networks more directly is via proteomic datasets, especially ones in which the abundance of the activated proteins is quantified (Sheng Wu and Radvanyi, 2010; Sachs et al., 2005; Brandi et al., 2022), though these tend to be more challenging experimentally.
@@ C.1 Links to datasets
In the following tables, we provide links to the different datasets we discussed in the main text. In Table 4, 5, 6 and 7 we report the links for pseudo-real, biology, neuroscience and Earth science datasets. We also keep an updated version of these lists of datasets at our github repo. The lists contain mainly datasets that have been used in causal discovery studies. Several are particularly interesting since they contain interventional data and can violate some common assumptions. Still, we advise thoughtful use in line with our recommendations, rather than applying them blindly. Also, the lists are by no means exhaustive and we do encourage researchers to explore new applications that are not listed here. While this section provides links to pseudo-real and real-world datasets, we would also like to share the synthetic generator Causally (Montagna et al., 2024) where it is natively possible to control the modelling assumptions (e.g., presence of latent confounders, unfaithful path, etc).
@@
(Cheng et al., 2024) CausalTime link to dataset
@@
We provide the links pointing to the original datasets. However, several of these datasets require some data preprocessing. In this paragraph, we elaborate on some datasets giving references to preprocessed versions that have been in causal discovery settings.
The two datasets from Replogle et al. (2022) have been processed and curated by Chevalley et al. (2022), and made accessible on CausalBench. The dataset from Frangieh et al. (2021) has been preprocessed by Lopez et al. (2022). The code is available online.
For the simulator SERGIO (Dibaeinia and Sinha, 2020), a modified version has been proposed by Hägele et al. (2023), making it possible to use custom graphs and generate (perfect) interventional data.
@@
DANDI Large collection of modern large neuroscience datasets, including optogenetics (DANDI) raw dataset link to database
(Dorkenwald et al., 2022) Drosophila connectome (Flywire) raw dataset (connectome) simulated Ca2+ activities
@@
(Teeters and Sommer, 2009) Spiking data from various sources (CRCNS) database
@@
(Shah et al., 2018) rs-fMRI data from the medial temporal lobe raw dataset
(Poldrack et al., 2015) Hippocampal rs-fMRI (MyConnectome project) raw dataset
@@
(Van Essen et al., 2013) rs-fMRI (Human Connectome Project) raw dataset
@@
(Wang et al., 2003) Task fMRI (star/plus experiment)
@@
Table 8: Links to Miscellaneous Real-World Datasets
Article Name Link
(Gamella et al., 2025) Causal Chambers link to simulator
(Wang and Körding, 2022) MOS 6502 microprocessor link to dataset
(Gentzel et al., 2019) Software Systems (PostgreSQL, JDK, Networking) link to datasets
(Huang et al., 2018) Archaeology data set
(Mooij et al., 2016) Cause-effect pairs (Tübingen pairs) link to dataset
(Byrne, 2013) Teacher’s burnout study link to dataset
(Shimizu et al., 2011) General Social Survey link to dataset